I have been helping my son with his homework and I myself need help w/ this problem.
Points X,Y, and Z are collinear, w/ point X between points Y and Z. If XY = 14.8 and YZ = 18.2, find XZ.

Thank you for the help.
Angela

Found 2 solutions by checkley75, Earlsdon:
Answer by checkley75(3666)

y--------x--------z
|-14.8---|
|-------18.2------|
18.2-14.8=3.4 for the distance between x&z.

Answer by Earlsdon(6294)

From the description of the problem, XZ = YZ-XY, so XZ = 18.2 - 14.8 = 3.4